Use the Henderson Hasselbalch equation to calculate the volume in mls of a) Tris acid and b) Tris base to use to make 250 ml of a 0.01M Tris buffer, pH 8.0, from 0.01M stock solutions. Give your answer to one place of decimals. pKa Tris = 8.3.
Answer #1

pH = pKa + log [Tris base] / [Tris acid]

8 = 8.3 + log [Tris base] / [Tris acid]

Moles of Tris base / Moles of Tris acid = 0.5011

Moles of Tris acid + Tris base = Molarity*V in L = 0.01*250/1000 = 2.5 mMoles

Tris base = 0.5011 Tris acid

0.5011 Moles of Tris acid + Moles of Tris acid = 2.5 mMoles

Moles of Tris acid = 0.001665

Moles of tris base = 0.0008345

Moles = Molarity *V in L

0.001665 = 0.01*V

V = 166.5 ml of Tris acid

0.0008345 = 0.01*V

V = 83.45 ml of Tris base